OpenStudy (anonymous):

Simplify the square root of 90

OpenStudy (anonymous):

\[\sqrt{90}\]

OpenStudy (anonymous):

3root10

OpenStudy (anonymous):

\[\sqrt{90}=\sqrt{9}\times \sqrt{10}=3\sqrt{10}\]

OpenStudy (anonymous):

9.49

OpenStudy (anonymous):

\[3\sqrt{10}\]

OpenStudy (anonymous):

yea \[3\sqrt{10}\]=9.49